# AMSNet: Netlist Dataset for AMS Circuits

### v1

@@ Abstract
Today’s analog/mixed-signal (AMS) integrated circuit (IC) designs demand substantial manual intervention. The advent of multimodal large language models (MLLMs) has unveiled significant potential across various fields, suggesting their applicability in streamlining large-scale AMS IC design as well. A bottleneck in employing MLLMs for automatic AMS circuit generation is the absence of a comprehensive dataset delineating the schematic-netlist relationship. We therefore design an automatic technique for converting schematics into netlists, and create dataset AMSNet, encompassing transistor-level schematics and corresponding SPICE format netlists. With a growing size, AMSNet can significantly facilitate exploration of MLLM applications in AMS circuit design. We have made an initial set of netlists public, and will make both our netlist generation tool and the full dataset available upon publishing of this paper.
@@ I Introduction
The recent emergence of MLLMs, such as GPT-4 and Gemini, presents great potential to bridge the gap of automatic AMS circuit design generation. There have been instances of MLLM-based circuit design works [3, 4, 5, 6, 7], yet comprehensive studies on the application of MLLMs in AMS circuits remain scarce. The challenge of MLLM-aided AMS circuit synthesis lies in knowledge acquisition and representation. A vast amount of data on AMS circuits, including schematic diagrams, netlists, and quantitative simulation results, is available online. However, this information often exists in only one of the necessary modals, complicating the extraction of cross-modal knowledge.
This issue does not significantly hinder the analysis and generation of digital circuits, as they typically require only modular-level netlists. Most likely, the training processes for modern MLLMs have included literature containing modular schematic diagrams and verbal explanations, enabling these models to identify functional modules for digital circuits naturally. On the contrary, AMS circuits are usually specified at the transistor level. Unfortunately, our experiments indicate that current MLLMs cannot extract transistor-level netlists with reasonable accuracy. Specifically, our preliminary experiments show that GPT-4 yields an accuracy of  90% in component detection, but only  45% accuracy in net detection.
@@
We first propose automatic netlist generation including component detection and net recognition. We then apply this to selected literature [10, 11] and form a preliminary version of dataset, titled AMSNet-1.0. Secondly, we conducted a series of experiments to explore the potential of GPT-4 in AMS circuit design. The results demonstrate GPT-4’s potential of providing design style suggestions given various application scenarios and specifications. Despite GPT-4’s limited ability to understand circuit topologies, AMSNet bridges the gap by establishing the correspondence between circuit schematics, netlists, and circuit functions as a source of knowledge. The combination of MLLMs and AMSNet realizes the automatic generation of AMS circuit topology.
Below section II presents AMSNet construction method and stats, and section III discusses ongoing and future work. AMSNet is expanded rapidly in terms of both database size and assoicated functions. An initial set of netlists is availble at our GitHub site https://ams-net.github.io/ams.net.github.io/. The entire dataset and source code to generate netlists from schematics will be made public upon publishing of this paper.
@@ II-C Netlist Generation
Figure 5: AMSNet-1.0 dataset summary on 894 schematics
@@ IV Conclusions
We have presented AMSNet, a database for AMS circuits. AMSNet is expanded rapidly in term of data size, circuit complexity, design sizing and specification, and assoicated AI functions. An initial dataset is availabe at Github site https://ams-net.github.io/ams.net.github.io/, and entire dataset and codes will be made public upon publishing of this paper.

### v2

@@ Abstract
Today’s analog/mixed-signal (AMS) integrated circuit (IC) designs demand substantial manual intervention. The advent of multimodal large language models (MLLMs) has unveiled significant potential across various fields, suggesting their applicability in streamlining large-scale AMS IC design as well. A bottleneck in employing MLLMs for automatic AMS circuit generation is the absence of a comprehensive dataset delineating the schematic-netlist relationship. We therefore design an automatic technique for converting schematics into netlists, and create dataset AMSNet, encompassing transistor-level schematics and corresponding SPICE format netlists. With a growing size, AMSNet can significantly facilitate exploration of MLLM applications in AMS circuit design. We have made the current version of database and associated generation tool public, both of which are expanded quickly.
@@ I Introduction
The recent emergence of MLLM presents great potential to bridge the gap of automatic AMS circuit design generation. There have been instances of MLLM-based digital circuit design [3, 4, 5, 6, 7], yet comprehensive studies on the application of MLLMs in AMS circuits remain scarce. The challenge of MLLM-aided AMS circuit synthesis lies in knowledge acquisition and representation. A vast amount of data on AMS circuits, including schematic diagrams, netlists, and quantitative simulation results, is available online. However, this information often exists in only one of the necessary modals, complicating the extraction of cross-modal knowledge.
This issue does not significantly hinder the analysis and generation of digital circuits, as they typically require only modular-level netlists. Most likely, the training processes for modern MLLMs have included literature containing modular schematic diagrams and verbal explanations, enabling these models to identify functional modules for digital circuits naturally. On the contrary, AMS circuits are usually specified at the transistor level. Unfortunately, our experiments indicate that current MLLMs cannot extract transistor-level netlists with reasonable accuracy. Specifically, our preliminary experiments show that GPT-4 yields an accuracy of about 90% in component detection, but only about45% accuracy in transistor-level net detection.
@@
We first propose automatic netlist generation including component detection and net recognition. We then apply this to selected literature [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and form a preliminary version of dataset, titled AMSNet-1.0. Secondly, we conducted a series of experiments to explore the potential of GPT-4 in AMS circuit design. The results demonstrate GPT-4’s potential of providing design style suggestions given various application scenarios and specifications. Despite GPT-4’s limited ability to understand circuit topologies, AMSNet bridges the gap by establishing the correspondence between circuit schematics, netlists, and circuit functions as a source of knowledge. The combination of MLLMs and AMSNet realizes the automatic generation of AMS circuit topology.
Below section II presents AMSNet construction method and stats, and section III discusses ongoing and future work. AMSNet is expanded rapidly in terms of both database size and assoicated functions. The latest dataset and generation tool are both publicly available at https://ams-net.github.io/.
@@ II-C Netlist Generation
Figure 5: AMSNet summary from 894 schematics
@@ IV Conclusions
We have presented AMSNet, a database for AMS circuits. AMSNet is expanded rapidly in term of database size, circuit complexity, and assoicated AI functions. All are available at https://ams-net.github.io/.